OpenStudy (anonymous):

Suppose that Y1,...,Yn is a random sample from Bernoulli(p). Find the distributions of Y(1)=min{y1,...,yn} and Y(n)=max{y1,...,yn}.

OpenStudy (anonymous):

order statistics... look it up

OpenStudy (anonymous):

Wow, so helpful.

OpenStudy (anonymous):

If I knew what do with it, I wouldn't be here....

OpenStudy (anonymous):

well, since it's bernoulli, it may simplify a bit. bernoulli is 1 with probability p. so the y's will either be 0's or 1's. Are the bernoulli's iid?

OpenStudy (anonymous):

and its' the \[y_i\] which are bernoulli, yeah? not the \[Y_i\]

OpenStudy (anonymous):

capital Y's throughout.

OpenStudy (anonymous):

what's iid?

OpenStudy (anonymous):

independent, identically distributed

OpenStudy (anonymous):

ah yes, they are. Bernoulli's are defined as independent. and their dist. is identical as stated in the problem.

OpenStudy (anonymous):

so Y_(1) can either be 0 or 1. it can be 1 iff all of the bernoulli's are 1, ow it's 0. P(Y_(1)=1)=P(all Y's are 1) = p^n => P(Y_(1)=0)=1-p^n

OpenStudy (anonymous):

for the max, you approach the other way round... P(Y_(n)=0) = P(all of the Y's are 0) = (1-p)^n P(Y_(n) = 1) = 1-(1-p)^n

OpenStudy (anonymous):

So they're both Bernoulli with a different p. For Y_(1) it's Bernoulli (p^n) and for Y_(n) it's Bernoulli [(1-p)^n]

OpenStudy (anonymous):

oops, i mean 1-(1-p)^n for Y_(n)

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

Kind of

OpenStudy (anonymous):

let's focus on the min...\[Y_{(1)}\] it's either going to be 0 or 1, right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so, how can it be 1? in other words, what has to happen in order for it to be 1?

OpenStudy (anonymous):

all would have to be 1

OpenStudy (anonymous):

thus the probability of that is p^n due to the independence

OpenStudy (anonymous):

right and since they're iid, =>\[ P(\text{all }Y_i \text{'s} = 1) = p \cdot p \cdot p \cdot (\cdots) \cdot p \,\,\text{(n \times)} = p^n\]

OpenStudy (anonymous):

but if the min is = 0?

OpenStudy (anonymous):

then since \[Y_{(1)}\] can only be 0 or 1, \[P(Y_{(1)}=0)=1-p^n\]

OpenStudy (anonymous):

thus, \[Y_{(1)}\sim\text{Bernoulli }(p^n)\]

OpenStudy (anonymous):

It's Bernoulli(p^n) because...?

OpenStudy (anonymous):

it can only be 0 or 1, right?

OpenStudy (anonymous):

yep and it's 1("success") with probability p^n

OpenStudy (anonymous):

got it.

OpenStudy (anonymous):

so now go over the max... and make sure you get that one too.

OpenStudy (anonymous):

good question... made me have to go into the deep recesses of my memory!

OpenStudy (anonymous):

it's Bernoulli with p =1-(1-p)^n

OpenStudy (anonymous):

yep, you understand how we got that?

OpenStudy (anonymous):

yeah, similar to the MIN, but start with Y=0

OpenStudy (anonymous):

you got it... all Y_i's are 0 in order for Y_(n) to be 0

OpenStudy (anonymous):

=>P( Y_(n)=0) = (1-p)^n thus P(Y_(n)=1) = 1 - (1-p)^n ("success")

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

have a look at this for other order stats...

OpenStudy (anonymous):

thanks for the question!